OpenStudy (anonymous):

Using u=π-x prove that integral from zero to π xf(sinx)dx=π/2*integral from zero to π f(sinx)dx

OpenStudy (anonymous):

\[\int\limits_{o}^{\pi}f(sinx)dx=(\pi/2)\int\limits_{0}^{\pi}f(sinx)dx\]

OpenStudy (anonymous):

Your initial question says \(\displaystyle \int_0^\pi x\cdot f(\sin(x)) \ dx\)

OpenStudy (anonymous):

Is that a typo?

OpenStudy (anonymous):

with the pi. I forgot to equate it